# Chirally-protected state manipulation by tuning one-dimensional statistics

### v1

@@ Abstract
$\hat{G}=\lim_{n\to\infty}\prod_{j=1}^{N_{\theta}}\hat{P}_{j},$ (8)
@@
For $L=2$ the zero energy subspace is non-degenerate and we calculate the (Abelian) Berry phase $\mathbf{G}=\pm 1$ exactly and find the nontrivial value of $-1$ for consecutive odd-integer coefficients in the Gauss sum $\sum_{k=1}^{N/2}k$, see [62]. We investigate the non-Abelian holonomies of larger systems by exact diagonalization [62, 64]. As an example, in Fig. 2(a), we present the holonomy matrix $\mathbf{G}$ in the chiral pair basis $|p(\mu)\rangle$ at $\theta=0$ for a system with $N=2$ and $L=6$ and $d_{0}=3$. The adiabatic evolution causes a significant non-trivial rotation of the zero energy subspace during the adiabatic evolution, which we indicate by ${\rho}_{ij}^{\nu}\left(\theta\right)=\langle\nu(\theta)|\hat{b}_{i}^{\dagger}\hat{b}_{j}^{\dagger}\hat{b}_{j}\hat{b}_{i}|\nu(\theta)\rangle$, an observable accessible in the corresponding experiments [31]. For example, we can take the initial two-body density $\rho_{ij}^{\nu}(\theta=0)$ of the basis state $|p(1)\rangle$ [Fig. 2(c)] and compare with ${\rho}_{ij}^{\nu}(\theta=2\pi)$, i.e., the two-body density corresponding to the adiabatically propagated state $\hat{G}|p(1)\rangle$. This yields a significant difference in the densities, see Fig. 2(d). Interestingly, for variations of parameters other than $\theta$ that preserve chiral symmetry, such as local variations in hopping strength $J$, we find that the connection is flat, meaning that $\hat{G}$ is not altered. Therefore, the results for the holonomy matrix $\mathbf{G}$ are robust against fluctuations in the adiabatic manipulation.
@@
This idea can be elevated to non-Abelian holonomies if we implement the site-dependent statistical parameter $\theta_{j}$. This leads to a generalized anyonic exchange algebra $a^{\dagger}_{j}a_{k}=\hat{a}_{i}\hat{a}_{j}^{\dagger}-e^{i\theta_{i,j}}\hat{a}_{j}^{\dagger}\hat{a}_{i}=\delta_{i,j}$ with $\theta_{i,j}=\theta_{j}$ for $i>j$, $\theta_{i,j}=-\theta_{i}$ for $i<j$, and $\theta_{i,i}=0$. For different exchange phases $\theta_{L}$ in the left and $\theta_{R}$ in the right part of the system [62], we generally find $\left[\mathbf{G}_{L},\mathbf{G}_{R}\right]\neq 0$, meaning the fluctuation-protected state manipulation depends on which statistical angle is altered first [62].
@@
Conclusions.— We have shown that chiral-symmetric number-preserving correlated hopping models with bosons accommodate at least $d_{0}$ degenerate zero energy states. The dimension of this space of states is robust against any parameter variations that preserve chiral symmetry, although the space itself sweeps through the chiral majority subspace. For the free anyon-Hubbard model, states in the zero energy chiral subspace can be prepared from experimentally-accessible initial states and detected by the characteristic checkerboard pattern in configuration space with current quantum gas microscopy techniques [31]. As the statistical angle is adiabatically tuned, the zero energy space picks up a non-Abelian holonomy for each cycle. This paradigm of non-Abelian state preparation could find further applications within the space of chiral symmetric correlated hopping models by tuning more than one cyclic parameter. As an example, we propose implementing this non-Abelian state preparation scheme within a spatially-inhomogeneous anyon-Hubbard model. More generally, we believe that Floquet-driven density-dependent Peierls phases offer a rich perspective for future exploration. Such models contain synthetic magnetic fluxes in few-body configuration space, and they provide an alternate path to understanding topological interactions in low dimensional systems.

### v2

@@ Abstract
$\hat{G}=\lim_{N_{\theta}\to\infty}\prod_{j=1}^{N_{\theta}}\hat{P}_{j},$ (8)
@@
For $L=2$ the zero energy subspace is non-degenerate and we calculate the (Abelian) Berry phase $\mathbf{G}=\pm 1$ exactly and find the nontrivial value of $-1$ for consecutive odd-integer coefficients in the Gauss sum $\sum_{k=1}^{N/2}k$, see [62]. We investigate the non-Abelian holonomies of larger systems by exact diagonalization [62, 64, 65]. As an example, in Fig. 2(a), we present the holonomy matrix $\mathbf{G}$ in the chiral pair basis $|p(\mu)\rangle$ at $\theta=0$ for a system with $N=2$ and $L=6$ and $d_{0}=3$. The adiabatic evolution causes a significant non-trivial rotation of the zero energy subspace during the adiabatic evolution, which we indicate by ${\rho}_{ij}^{\nu}\left(\theta\right)=\langle\nu(\theta)|\hat{b}_{i}^{\dagger}\hat{b}_{j}^{\dagger}\hat{b}_{j}\hat{b}_{i}|\nu(\theta)\rangle$, an observable accessible in the corresponding experiments [31]. For example, we can take the initial two-body density $\rho_{ij}^{\nu}(\theta=0)$ of the basis state $|p(1)\rangle$ [Fig. 2(c)] and compare with ${\rho}_{ij}^{\nu}(\theta=2\pi)$, i.e., the two-body density corresponding to the adiabatically propagated state $\hat{G}|p(1)\rangle$. This yields a significant difference in the densities, see Fig. 2(d). Interestingly, for variations of parameters other than $\theta$ that preserve chiral symmetry, such as local variations in hopping strength $J$, we find that the connection is flat, meaning that $\hat{G}$ is not altered. Therefore, the results for the holonomy matrix $\mathbf{G}$ are robust against fluctuations in the adiabatic manipulation.
@@
This idea can be elevated to non-Abelian holonomies if we implement the site-dependent statistical parameter $\theta_{j}$. This leads to a generalized anyonic exchange algebra $a^{\dagger}_{j}a_{k}=\hat{a}_{i}\hat{a}_{j}^{\dagger}-e^{i\theta_{i,j}}\hat{a}_{j}^{\dagger}\hat{a}_{i}=\delta_{i,j}$ with $\theta_{i,j}=\theta_{j}$ for $i>j$, $\theta_{i,j}=-\theta_{i}$ for $i<j$, and $\theta_{i,i}=0$. For different exchange phases $\theta_{L}$ in the left and $\theta_{R}$ in the right part of the system [62, 66], we generally find $\left[\mathbf{G}_{L},\mathbf{G}_{R}\right]\neq 0$, meaning the fluctuation-protected state manipulation depends on which statistical angle is altered first [62].
@@
Conclusions.— We have shown that chiral-symmetric number-preserving correlated hopping models with bosons accommodate at least $d_{0}$ degenerate zero energy states. The dimension of this space of states is robust against any parameter variations that preserve chiral symmetry, although the space itself sweeps through the chiral majority subspace. For the free anyon-Hubbard model, states in the zero energy chiral subspace can be prepared from experimentally-accessible initial states and detected by the characteristic checkerboard pattern in configuration space with current quantum gas microscopy techniques [31]. As the statistical angle is adiabatically tuned, the zero energy space picks up a non-Abelian holonomy for each cycle. This paradigm of non-Abelian state preparation could find further applications within the space of chiral symmetric correlated hopping models by tuning more than one cyclic parameter. As an example, we propose implementing this non-Abelian state preparation scheme within a spatially-inhomogeneous anyon-Hubbard model. More generally, we believe that Floquet-driven density-dependent Peierls phases offer a rich perspective for future exploration. Such models contain synthetic magnetic fluxes in few-body configuration space, and they provide an alternate path to understanding topological interactions in low dimensional systems. We propose a generalized anyon-Hubbard model with inhomogeneous statistical angles as one example of wider class of bosonic Hamiltonians with density-dependent Peierls phases, possibly without anyonic interpretation, but relevant to experiments and topological control via adiabatic holonomies.